ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

10. Partial Differential Equations

1991 MS Classification: 35B27, 35B40 Chechkin, Gregory A., Moscow State University, Russia On Behavior of Perforated Bodies with Rapidly Oscillating Surfaces. We consider boundary-value problem for the Poisson equation in a perforated domain with rapidly oscillating exterior boundary in the presence of a small dissipative term in the boundary conditions. We study this problem assuming that both perforation and boundary oscillation are locally periodic and their structures are consistent. We prove the homogenization theorem and estimate the rate of convergence. I hope that it is not too late. Thank you very much in advance. I have an additional problem. My wife, Chechkina Tatiana P. , is a mathematician too. She wants to attend this conference with me. I would like to ask you to send me the invitation for my wife by air mail. In my application form I have written that we will take part at ICM98 together, but I have not received the invitation for my wife yet. She needs this invitation for visa. Short Curriculum Vitae of Chechkina Tatiana P. : Assistant Professor Moscow State Textile Academy, Moscow If it is possible send me please the invitation for my wife using my address: Differential Equation department Facuty of Mechanics and Mathematics Moscow State University Moscow 119899, Russia We need this invitaton as soon as possible. Sorry for any inconvenience it will may cause. Thank you. Hope hearing from you soon. Best wishes. Sincerely yours, Gregry A. Chechkin. —– End of forwarded message from Gregory A. Chechkin —-